# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ 4.1 Consumption Smoothing
I consider a linear fixed-effects model to test consumption smoothing. [^51] Another empirical specification used to test consumption smoothing is the first-difference model, including the aggregate consumption to capture macroeconomic shocks in the economy (Mace 1991). This alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. However, the THBS households do not include entire households in the target economy—Tsukishima. This may disturb the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance. [^52] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption-smoothing literature. Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences (TWFE-DID) setting yet. In fact, the bias-corrected estimator for the TWFE-DID model with continuous treatment for multiple periods is underdeveloped (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review). The cluster-robust variance-covariance matrix estimator is used to manage heteroskedasticity and serial dependency (Arellano 1987).
@@ Results
(1) Semi-monthly Panels (2) Adjusted Monthly Panels
@@
Notes: This table shows the results of equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed. Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumption. An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant. This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to manage those shocks to smooth food consumption. The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant. This may be understandable because those expenses include family travel and theatergoing. [^53] The workers could walk to the factories in Tsukishima, so the costs of commuting were not usually documented in the budget books. Similarly, the estimate for gifts shows a large amount ($0.523$), reflecting that the expenses on the others were optional. The results for the other subcategories are statistically insignificant or unclear. The estimates for furniture and clothes are weakly statistically significant. Although not conclusive, this may reflect that both categories were luxury categories. The estimate for the education category is small but also weakly statistically significant ($0.101$). This may be because it includes an allowance for children, which could be easily decreased. [^54] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University. He left important retrospectives on the survey in his works (Online Appendix B.1). According to his book, since there were no nurseries, children in these working-class households regularly received a small allowance to play outside while their parents were working (Miyoshi 1989, pp. 383–384; 386; 387).
@@
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed over a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption was $0.433$, which is approximately $1.2$ times greater than that from the semi-monthly series. One potential explanation for this difference is that the available sources of temporary income might have been limited, and those sources were not useful for compensating for adverse income shocks that lasted for at least one month. The most noticeable example is the food category, which comprised the largest portion of the expenditure (Panel A of Table 3). The estimate is $0.293$, more than twice the estimate for the semi-monthly series ($0.136$). This implies that the households’ food expenditures were likely to be insured for a few weeks but not over one month. I will investigate the mechanism underlying this behavior in the following subsections, and credit purchases may be used to smooth food consumption in the short run. Similarly, the expenses for entertainment and transportation were much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series. The estimate for gifts is slightly smaller but in a similar range. In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The expenses on those unnecessary categories might particularly be cut in response to relatively persistent shocks.
@@
In contrast, a few subcategories were not sensitive to the idiosyncratic shocks over different time-series frequencies. For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect the fact that the rent and utility payments were relatively fixed. [^55] For example, rent payment was usually fixed and paid once per month. Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series. The medical expenses included the fee for bathing (yuya). People took baths every few days at that time, which was less likely to be cut, although hairdressing for women (kamiyui) in the same category was likely to be cut. The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series. While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, such as unanticipated losses of household goods, clothes, and stationery, including writing paper. [^56] An alternative explanation for the clothes category relates to seasonality. The workers had bought their work clothes for the new season (e.g., the OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)). As those clothes were necessary items, those seasonal expenses may attenuate the estimate. This may systematically attenuate the estimates for the specific categories in the adjusted monthly series. [^57] Generally, the panels with short-run time bins were more likely to be influenced by unexpected consumption shocks, which attenuated income elasticity (Nelson 1994). If this mechanism worked, subcategories that were sensitive (insensitive) to unexpected consumption shocks should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels in a large part of subcategories, suggesting that households faced more difficulties with smoothing their consumption in the long run.
@@ 4.2 Risk-coping Mechanisms
The preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, became smaller when I focused on the semi-monthly frequency. Specifically, food expenditure was likely to be insured within at least half a month. The workers might have coped with their shorter-run shocks relatively well but could not manage longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables in the risk-coping strategies.
@@
To test the roles of risk-coping strategies, I employed the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I used the head’s income instead of the disposable income in Equation 2 to avoid the potential endogeneity. [^58] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results were largely unchanged if I used the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I preferred to use the conservative setting to manage potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. Regarding $r_{i,t}$, I considered several variables to overview various risk-coping strategies. These included the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable was defined as the difference between income and expenses. The estimates of $\kappa$ were therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also considered the earnings of the family members (except for the head) and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales might alleviate idiosyncratic shocks as alternative risk-mitigation strategies (Udry 1995; Blundell et al. 2016). [^59] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focused on the households living in Tsukishima throughout the sample period, I did not need to address the ex-ante risk-mitigation strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
(1) Semi-monthly Panels (2) Adj. Monthly Panels
Head’s earning Obs. Head’s earning Obs.
@@
Panel A: Savings, insurance, borrowing, and gifts (yen)
@@
Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
@@
Notes:
This table shows the results of equation 2. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
Table 5 presents the results. Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively. Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of small assets.
@@
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the adjusted monthly series is similar ($0.035$), albeit statistically insignificant (Column (2)). This is consistent with the historical fact that many workers had savings accounts in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude was small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
@@
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in this category was the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating for the idiosyncratic shocks. For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which is independent of the worker’s earnings. [^60] The OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the income from mujin measured in the budget books was less than one yen, which is not the case on average.
@@
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshops were a popular lending institution for low-income households. Although THBS households occasionally used pawnshops, this was not frequent. [^61] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This is understandable because skilled factory workers were not classified as low-income working class at that time.
@@
The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column (1). The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is the largest in all the categories. According to the THBS budget books, the credit purchases were mainly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). As rice—a staple food—is usually an expensive item, credit purchases might have played an important role among these households. In addition, credit purchase was occasionally used for the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While these results suggest that credit purchase was the primary risk-mitigation strategy, Column (2) shows that credit was no longer effective in the long term. This is reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Given the estimated magnitude ($0.13$), however, it is noteworthy that the credit could only partially compensate for the idiosyncratic shocks.
@@
The estimates for the gifts category are negative and statistically significant in Column (1). The budget books show that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be approximately $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns of the first row are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that the wives regularly engaged in side jobs rather than working temporarily to compensate for the income loss. In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^62] The OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919). In the second row, I show that selling miscellaneous assets such as newspapers, bottles, and clocks was not an effective risk-mitigation strategy.
To summarize, the evidence suggests that credit purchases mitigated idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were available for the short term —at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions. This supports the findings on partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding—which is not explicitly described in the historical sources—is that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.

### v14

@@ 4.1 Consumption Smoothing
I use a linear fixed-effects model to test consumption smoothing. Online Appendix C.1 presents the conceptual framework and the derivation of my empirical specification. For household $i$ at time $t$, the reduced form equation is
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures time-invariant unobservables such as permanent income and consumption preferences. The time fixed effect accounts for unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ represents income elasticity, ranging from zero under perfect insurance to one in the absence of insurance. Since identification relies on both adverse and favorable idiosyncratic income shocks, the estimated elasticity reflects the overall response to these shocks. A cluster-robust variance-covariance matrix estimator is used to address heteroskedasticity and serial correlation (Arellano 1987).
Another specification commonly used in the literature is the first-difference model, which uses aggregate consumption to capture macroeconomic shocks (Mace 1991). However, the aggregate measure may not accurately reflect macroeconomic shocks, particularly when the number of households is relatively small. To be conservative, I therefore employ the two-way fixed-effects model, which systematically controls for macroeconomic shocks through the time fixed effect (Cochrane 1991; Ravallion and Chaudhuri 1997).
@@ Results
(1) Semi-monthly panels (2) Monthly panels
@@
Notes: This table presents the results of equation 1. It reports regressions of 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, the household fixed effect, and the time fixed effect. Columns 1 and 2 show the results for the adjusted semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are shown in the columns labeled “Coef.” The estimated elasticities reflect the overall responses of households to both adverse and favorable idiosyncratic income shocks.
@@
Column 1 of Table 4 presents the results from the semi-monthly panel dataset. The estimated coefficient for total consumption indicates that a one percentage point decrease in disposable income reduces total expenditure by $0.358$ percentage points, suggesting that factory-worker households could not fully smooth their consumption. An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant. This suggests that while households reduced food expenses in response to adverse income shocks, they still had some means to smooth food consumption. The estimates for entertainment and transportation expenses ($0.307$ and $0.278$) are also statistically significant. This is understandable because these expenses include family travel and theater visits. [^39] On this small island, workers could walk to the factories, so commuting costs were typically not recorded in the budget books. Similarly, the larger estimate for gifts ($0.523$) reflects that these expenses were discretionary. The results for other subcategories are statistically insignificant or ambiguous. Estimates for furniture and clothing are weakly statistically significant, which may reflect that both are luxury categories. The estimate for the education category is small but also weakly statistically significant ($0.101$), as it includes an allowance for children that can be easily reduced. [^40] Toyotaro Miyoshi, who served as a research assistant at the Tsukishima Survey, provided important retrospective accounts (Online Appendix B.1). According to his book, children in these factory-worker households regularly received a small allowance to play outside while their parents worked, since no nurseries were available (Miyoshi 1989, pp. 383–384; 386; 387).
@@
Column 2 reports the results for the monthly panel dataset examining whether income elasticities change over a longer period. As expected, the monthly series yields larger estimates than the semi-monthly series. The estimate for total consumption is $0.433$, approximately $1.2$ times higher than that from the semi-monthly series. One possible explanation is that the sources of temporary income are limited and insufficient to offset adverse income shocks lasting at least one month. The most notable example is the food category, which shows more than twice the estimate ($0.293$) compared with the semi-monthly series. This implies that households’ food expenditures could be insured for a few weeks but not over a full month. Similarly, expenses for entertainment and transportation are much more sensitive ($0.716$ and $0.605$, respectively) to shocks in the monthly series. The estimate for gifts is slightly lower but remains in a similar range. Expenses on these discretionary categories are particularly responsive to relatively persistent shocks.
@@
By contrast, some subcategories are not sensitive to idiosyncratic shocks across different time-series frequencies. For example, estimates for housing, utilities, and medical expenses are not statistically significant in either the semi-monthly or monthly series. This may reflect the fact that rent and utility payments are relatively fixed. [^41] Rent is usually fixed and paid once per month. Online Appendix C.2 provides quantitative evidence that the housing category should be analyzed using the monthly series rather than the semi-monthly series. Medical expenses include fees for bathing (yuya). People bathe every few days, a habit unlikely to change, although the frequency with which women opt for hairdressing services(kamiyui) in the same category can reduce. The estimates for furniture, clothing, and education are statistically insignificant in the monthly series. While this may partly reflect efficiency issues, these expenses are generally insensitive to unexpected demand shocks such as unanticipated losses of household goods, clothing, or stationery, including writing paper. [^42] An alternative explanation for the clothing category relates to seasonality. Workers bought new work clothes for the season (OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)). As these clothes were necessary, seasonal purchases attenuate the estimates. This systematically reduces the estimates for specific categories in the monthly series. [^43] Panels with short-run time bins are more likely to be influenced by unexpected consumption shocks, which attenuate income elasticity (Nelson 1994). When this mechanism operates, subcategories sensitive (insensitive) to unexpected consumption shocks are more (less) attenuated. In my results, however, estimates for the monthly panels are higher than those for the semi-monthly panels for many subcategories, suggesting that households faced greater difficulties smoothing consumption over the longer term.
@@ 4.2 Risk-coping Mechanisms
The preceding results suggest that workers were able to cope with short-term shocks to some extent. In this section, I examine the mechanisms underlying their consumption-smoothing behavior using information on risk-coping strategies.
@@
To test the roles of risk-coping strategies, I employ the empirical specification proposed by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is
@@
where $\tilde{y}_{i,t}$ is the head’s earnings, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I use the head’s income instead of total disposable income in Equation 2 to avoid potential endogeneity. [^44] For example, total household income is the outcome of intra-household responses to shocks, such as labor supply adjustments (Asdrubali et al. 2020). The results are largely unchanged if disposable income is used instead of the head’s income (not reported). This is consistent with the absence of labor supply adjustments among THBS households (Table 5). Nevertheless, I use the conservative approach to reduce potential biases and maintain comparability with the structural equation in Section 4.3. Regarding $r_{i,t}$, I consider several variables to capture different risk-coping strategies, including net income from savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as income minus expenses. The estimates of $\kappa$ are expected to be negative, since net income is typically negative in good times and positive in unfavorable times. I also consider earnings of other family members (excluding the head) and income from sales of miscellaneous goods (Udry 1995; Blundell et al. 2016). [^45] Rosenzweig and Stark (1989) show that marriage combined with migration in South Indian villages reduced variability in household food consumption. Since the THBS tracked households living in Tsukishima throughout the sample period, ex-ante risk mitigation via migration is not relevant here. Estimated coefficients for sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income declines.
@@ Results
(1) Semi-monthly panels (2) Monthly panels
Head’s earnings Obs. Head’s earnings Obs.
@@
Panel A: Net savings, insurance, borrowing, and gifts (income - expenditure, yen)
@@
Other members’ earnings 0.001 [0.010] 289 0.036 [0.033] 124
@@
Notes: This table reports the results of equation 2. All regressions include the family size control, household fixed effects, and time fixed effects. Columns 1 and 2 present estimates for the adjusted semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are shown in the “Coef.” columns. Panel A: Each net income variable is defined as income minus expenditure. “Net savings”: withdrawals minus deposits to savings. “Net insurance”: insurance payouts minus insurance expenses. “Net borrowing”: borrowing minus debt repayments. This category includes net income from both pawnshops and other lending institutions. “Net credit purchase”: credit purchases minus credit repayments. “Net gifts”: the total received gifts, both pecuniary and non-pecuniary, minus payments for gifts. Panel B: “Other members’ earnings”: total income earned by all household members except the head. “Sales of miscellaneous assets”: include the sale of daily items such as newspapers and empty bottles. All dependent variables are measured in yen.
@@
Table 5 presents the results. Columns 1 and 2 summarize the findings for the semi-monthly and monthly panel datasets, respectively. Panel A reports the results for the net income variables, while Panel B reports those for labor supply adjustments and sales of small assets.
@@
The estimated coefficient for net savings is negative and statistically significant in Column 1. This suggests that a one yen decrease (increase) in the head’s earnings increased (decreased) net savings by $0.043$ yen. The estimate from the monthly series is similar but statistically insignificant (Column 2). While this reflects that many workers had savings accounts in postal savings and savings banks (Section 2.3), the magnitude was small, indicating that withdrawals from savings did not fully compensate for earnings losses.
@@
Estimates for the insurance category are near zero and statistically insignificant in both columns. The budget books indicate that the primary source was mutual loan associations (mujin). However, these associations’ rules were unsuitable for offsetting idiosyncratic shocks. For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which was independent of the worker’s earnings. [^46] OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919). The bidding rule could support a specific member if all members have a common belief in aid, but most income from mujin recorded in the budget books was less than one yen, which is not representative on average.
@@
For net borrowing, estimated coefficients are close to zero and statistically insignificant in both columns. Although THBS households occasionally used pawnshops, these instances were infrequent. [^47] OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This is understandable because skilled factory workers were not classified as low-income, who were the primary users of pawnshops (Section 2.3).
@@
The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column 1. This suggests that a one yen decrease (increase) in the head’s earnings increased (decreased) net income from credit purchases by $0.13$ yen. This magnitude is the largest among all categories. According to the budget books, credit purchases were primarily used to buy rice, fish, vegetables, and fuel (firewood and charcoal). As rice—a staple food—was usually expensive, credit purchases played an important role for these households. Credit purchases were also occasionally used for medical expenses such as doctor visits and pharmacy purchases, consistent with the baseline results for medical expenses (Column 1 in Table 4). While these results indicate that credit purchases were the primary risk-mitigation strategy in the short term, Column 2 shows that credit ceases to be effective over the longer term. This is reasonable, as borrowers typically made monthly or semi-monthly payments and did not borrow for periods longer than a month (Section 5).
@@
Estimates for the gifts category are negative and statistically significant in Column 1. The budget books show that the gifts include seasonal gifts, celebration money, and return gifts. Most of these gifts were not aid from kinship networks, including rural relatives, but rather monetary and material exchanges with acquaintances based on a culture of gift-giving. According to an official report from Tokyo City, average monthly gifts accounted for approximately 33% of total income (Social Affairs Division 1925, pp. 58-59). Similarly, this share is approximately 44% in my monthly panel dataset. Even if mutual aid from personal networks is included in this small figure, it likely provided insufficient support for households facing economic hardship. Accordingly, the estimate ceases to be statistically significant in the monthly series (Column 2) and becomes negligible when focusing on the impact of adverse income shocks (Section 4.3).
@@
Finally, Panel B tests the roles of intra-household labor supply adjustments and sales of miscellaneous assets. The estimated coefficients in both columns of the first row are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that wives regularly engaged in side jobs such as sewing, rather than taking temporary work to compensate for income losses. However, their earnings were too small to replace the lost income from their breadwinning husbands. [^48] OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919). In the second row, sales of miscellaneous assets such as newspapers, bottles, and clocks are shown to be ineffective as a risk-mitigation strategy.
Heterogeneity
The residential area of Tsukishima spanned less than $1.5$ kilometers, with a width of $500$ meters (Chūō Ward 1994). Within this narrow area, workers had easy access to shops in different lots (Department of Health, Ministry of the Interior 1923a, pp. 40–46). Nevertheless, I conduct several additional tests to explore heterogeneous responses using maps of retailers and pawnshops documented in the report.
First, I calculate the density of retail shops for daily necessities in each lot and include it in the regression for credit purchases. The estimate from this expanded regression is largely consistent with my baseline estimate (Column 1 in Panel A of Table 5). As an alternative test, I add an interaction term between the head’s earnings and the density of retailers into Equation 2. The interaction effect is approximately zero and remains unchanged when using an alternative definition of density. These results indicate that shop accessibility was homogeneous for the workers. Online Appendix C.3 provides detailed summaries of these analyses.
Second, I examine the role of local pawnshops separately from the net borrowing regression in Table 5. In European countries and the U.S., religious communities played an important role in providing social insurance (Ferrara and Testa 2023; Ager and Ciccone 2018). Although some philanthropists assisted the poor before government-managed social welfare became available, few social facilities run by religious communities existed in Japan (Yoshida 2004). Unlike in European countries (Murhem 2016), pawnshops in Japan were private, profit-oriented enterprises until public pawnshops with social functions were established in 1927 under the Public Pawnbroking Act. Despite this, local pawnshops may have had informal ties with workers, as they served the same residents of Tsukishima.
When I use temporary income from pawnshops as the dependent variable in Equation 2, the estimated coefficient on the head’s earnings is close to zero and statistically insignificant. [^49] The cooperative society is another local risk-coping institution, though it is rarely used (Section 2.3). Using total purchases from the cooperatives as a dependent variable in Equation 2 yields a coefficient on the head’s earnings that is nearly zero and statistically insignificant (Online Appendix C.4). This result remains unchanged after including the density of pawnshops in each lot in the same regression. I then estimate an expanded regression of Equation 2 for temporary income from pawnshops, which includes an interaction term between the head’s earnings and the density of pawnshops. Although the interaction effect is negative, it is small and statistically insignificant. A similar result is obtained when using an alternative definition of the density variable. These findings suggest that the use of pawnshops was independent of idiosyncratic shocks. Online Appendix C.4 summarizes the results.